How many nickles would you trade for 250yen? $1 = 150yen
1
Expert's answer
2012-10-04T08:42:27-0400
Taking into account that $1 = 20 nickles, we have

$1 = 150yen ==> 1yen = $1/150 ==> 250yen = $250/150 ==> 250yen = 250*20/150 nickles = 33 + 1/3 nickles.